16. Solve in |R:

(a) 2(3y-2)+3(4-3y)>12-(5y+7)

(b) 2(k+2)+3(2k-3)<or=2


17.(a) if x=1-3t and y=2t-1.

Find the value of t such that x=y.

(b) solve for real value of x the equation,

(m+1)x²-(m+3)x+3-m=0.

When m=-1.


18.Given the parametric equations:

x=t-1/t

y=2(t+2)/t²

Find the coordinates of the point where,

(a) t=-2

(b) t=-1


19.solve in |R:

(a) 4(m+5)-6(2m+3)=3(m+14)-2(5-m)+9

(b) 3(k-1)-4(k+6)=10(k+5)



1
Expert's answer
2022-02-10T13:11:04-0500

16.

a)

"2(3y-2)+3(4-3y)>12-(5y+7)"

"6y-4+12-9y-12+5y+7>0"

"2y+3>0"

"y>-3\/2"

(b)


"2(k+2)+3(2k-3)\le2"

"2k+4+6k-9\le 2"

"8k\le7"

"k\le7/8"

17.

(a)


"x=y=>1-3t =2t-1."

"5t=2"

"t=2\/5"

(b)


"(-1+1)x^2-(-1+3)x+3-(-1)=0"

"-2x+4=0"

"2x=4"

"x=2"

18.

(a)

"x=-2-1\/(-2)=-3\/2"

"y=2(-2+2)\/(-2)^2=0"

Point "(-3\/2, 0)"

(b)

"x=-1-1\/(-1)=0"

"y=2(-1+2)\/(-1)^2=2"

Point "(0, 2)"


19.

(a)


"4(m+5)-6(2m+3)=3(m+14)-2(5-m)+9"

"4m+20-12m-18=3m+42-10+2m+9"

"-8m+2=5m+41"

"13m=-39"

"m=-3"

(b)


"3(k-1)-4(k+6)=10(k+5)"

"3k-3-4k-24=10k+50"

"-k-27=10k+50"

"11k=-77"

"k=-7"